# Navigating the Labyrinth: Path-Sensitive Unit Test Generation with Large Language Models

### v1

@@ Abstract
11footnotetext: Corresponding author.
@@ I Introduction
To address this, we design and implement JUnitGenie, a path-sensitive framework that guides LLMs in generating high-coverage test cases with minimal cognitive overhead, i.e., by supplying only the essential context needed while filtering out irrelevant information. Specifically, we first build a code knowledge base (CKB) by extracting contextual knowledge (i.e., type information, control flow, and data dependencies) from a Java project under testing. We then traverse each method one control-flow path at a time, retrieve path-specific context from the CKB, and distill it into explicit instructions to guide test generation for that specific path. JUnitGenie applies multi-round refinement to resolve compilation and runtime errors, iteratively generating test cases that are both syntactically correct and executable at runtime without errors.
@@
  • We conduct an extensive evaluation to demonstrate the effectiveness of JUnitGenie: (i) it significantly outperforms state-of-the-art tools across real-world Java projects; (ii) ablation studies confirm the contribution of knowledge distilllation and test case refinement; (iii) the approach generalizes well across popular foundation models; (iv) it successfully uncovers real bugs in production code.
@@ III-A Context Knowledge Extraction
2) Control-flow information: We then use SootUp [29], which offers precise features for code analysis. In particular, it supports 3-address code intermediate representation Jimple and accurate call-graph construction algorithm Spark [30]. On top of SootUp, JUnitGenie decomposes Java programs into Jimple and decomposes expressions into CFG paths, modeling method calls as atomic nodes and logical operators as branching edges. Each path has distinct guard conditions, enabling input synthesis for diverse execution scenarios. The CFG then captures feasible execution paths, helping identify conditional branches and guiding diverse test-input generation. When extracting paths from a CFG, if a loop is encountered (i.e., a node reappears), we only keep the first iteration to preserve the loop’s logic, and skip subsequent visits within the same path to prevent infinite path extraction.
3) Data-flow dependencies: We again adopted SootUp to extract data-flow information such as variable definitions, field accesses, and dependent method calls to capture how values may influence branching conditions. Specifically, dependent methods are those whose return values affect subsequent control decisions or outputs. For example, as shown in Fig. 1, the return value of contains() (i.e., line 12) determines whether the method proceeds to the following return true statement at line 13. In addition, JUnitGenie also considers dependent variables, including parameters, local variables, and static fields whose values are referenced in conditional statements within the focal method. For instance, the static string variable VOWELS is used in isVowel() to determine whether a character belongs to the set of vowels, thereby influencing the outcome of the conditional branch at line 16. Understanding this context knowledge and their data flow enables LLM to assign appropriate values to variables when constructing test inputs. Together, these dependencies provide a detailed view of how data influences control flow, allowing JUnitGenie to generate inputs that satisfy path-specific conditions and fully exercise the focal method’s behavior.
@@ IV-A Dataset
To evaluate the effectiveness of JUnitGenie, we use real-world Java projects from Defects4J [36], a widely adopted benchmark for automated testing research. We use the latest versions of these projects to reflect modern development practices. In addition, we observe Defects4J includes 17 Java projects with varying JDK requirements. To ensure a consistent and controlled experimental environment, we standardize on JDK 8 as the target platform, given its continued prevalence in open-source libraries and enterprise systems [37, 38, 39]. Based on this criterion, we exclude four projects (JFreeChart, Commons-CLI, Jsoup, and Math) that require JDK versions beyond 8 and cannot be compiled under our configuration. Additionally, exclude two projects (Mockito and Closure-Compiler) that are not Maven-based and therefore cannot be built using our toolchain. As a result, 10 Java projects remain and are used as our experimental dataset, as shown in Table I.
Within these ten projects, not all methods are equally meaningful for evaluating test generation tools. Specifically, methods with trivial logic (e.g., no branching or dependencies) can be easily covered by naive approaches and fail to showcase the strengths of advanced tools. To ensure rigorous evaluation, we focus on methods that are both challenging and representative of real-world complexity. Specifically, we select focal methods that meet two criteria: 1) contain at least one branching condition; and 2) have data dependencies on variables or methods defined outside its own body. These filters yield 8,367 candidate methods across the ten projects. However, generating tests for all would incur significant computational and token costs. To balance cost with statistical significance, we randomly sample [^2] The sample size is determined based on a confidence level at 95% and a confidence interval at 5 (https://www.surveysystem.com/sscalc.htm). a subset of methods from each project, as summarized in Table I. Hence, the overall dataset contains 2,258 focal methods spread across 762 classes.
@@ IV-B RQ1-Effectiveness
Our first research question concerns the effectiveness of JUnitGenie in test case generation. To this end, evaluate JUnitGenie by measuring its ability to produce valid tests and comparing its performance with existing SOTA tools. Here, we consider a test case to be valid if (1) the generated code snippet can be successfully compiled and (2) it does not throw an exception before the execution point of the focal method. The first condition ensures that the test case is syntactically correct. The second condition guarantees that the execution context of the focal method under test is properly constructed so that the method can be successfully reached during execution.
@@
In addition, Table II reports the number of valid test cases generated by each tool, showing that heuristic-based approaches yield fewer valid tests than LLM-based ones. Specifically, EvoSuite and Randoop generate 3,232 and 4,018 valid tests, while LLM-based tools ChatTester, CoverUp, HITS, and JUnitGenie produce 6,324, 7,971, 8,351, and 14,190 valid tests, respectively. These results highlight the superior effectiveness of JUnitGenie in generating executable test cases.
Fig. 4(a) further summarizes the distribution of valid test cases per project for each tool. Overall, LLM-based tools consistently produce more valid test cases than heuristic-based approaches. Specifically, the median number of valid test cases per project is 323 for EvoSuite, 415 for Randoop, 260 for ChatTester, 332 for CoverUp, 432 for HITS, and 1,355 for JUnitGenie, with JUnitGenie also exhibiting the highest upper quartile at 1,633. Like most other SOTA approaches, JUnitGenie may fail on focal methods with complex control or data dependencies. Nonetheless, JUnitGenie is still capable of generating more valid test cases than existing tools, demonstrating its effectiveness in leveraging distilled context across diverse codebases.
Moreover, we compare the branch and line coverage achieved by each tool across the ten Java projects, as shown in Table III. JUnitGenie consistently achieves the highest average coverage, with 56.86% for branch and 61.45% for line coverage, outperforming both heuristic and LLM-based baselines. It delivers particularly strong results on complex projects such as Lang (87.10% branch coverage, 92.41% line coverage) and CSV (76.57% branch coverage, 85.98% line coverage), demonstrating its ability to generate tests that exercise diverse and deep execution paths. Although tools like EvoSuite and Randoop perform well on a few individual projects, their overall average coverage remains substantially lower. Among LLM-based methods, HITS achieves the second-best performance with 32.13% branch and 36.45% line coverage, falling notably behind JUnitGenie. These results further validate the effectiveness of JUnitGenie in producing high-coverage test suites across a range of real-world Java applications. We then break down the comparative results as follows:
@@ IV-C RQ2-Knowledge Acquisition and Utilization
While JUnitGenie demonstrates strong performance in generating high-coverage test cases, the specific contributions of its core components remain unclear. To answer this question, an ablation study is conducted to assess the impact of three key steps: Context Knowledge Extraction (CKE), Context Knowledge Distillation (CKD), and Test Case Refinement (TCR). The study aims to clarify how each component contributes to the overall effectiveness of the test generation process.
Here, we design four ablation baselines to assess the contribution of each component in JUnitGenie. The most basic variant (w/o CKD+CKE+TCR) removes all modules, providing only the focal method and its signature. The second (w/o CKD+CKE) retains only test case refinement, using error messages to repair tests without any contextual input. The third (w/o CKD) includes both context extraction and refinement but omits distillation, presenting raw context without structured guidance. The full JUnitGenie configuration includes all three components. All baselines use GPT-4o-Mini and are evaluated on 2,258 focal methods from ten real-world Java projects.
TABLE V: Comparison of branch and line coverage using basic prompting (Basic) vs. JUnitGenie across four LLMs.
@@
    GPT4o-mini CodeLlama DeepSeek-Coder Qwen2.5-Coder     GPT4o-mini CodeLlama DeepSeek-Coder Qwen2.5-Coder
@@
Adding test case refinement (w/o CKD+CKE) improves coverage by 12–15%, as it repairs syntactically invalid yet semantically meaningful tests through iterative refinement of compiler error messages. As illustrated by the red line in Fig. 6, refinement raises the executable test rate from 33.49% to 65.81% over five iterations. Other variants show substantial gains over five rounds, with JUnitGenie improving from 45.60% to 69.88%, and w/o (CKD+CKE) from 20.32% to 63.44%. These trends highlight the critical role of iterative refinement in recovering test cases, especially when combined with contextual knowledge. It is also worth noting that JUnitGenie, which incorporates both contextual knowledge and refinement, achieves the highest improvement, reaching 69.88% by round five. In comparison, w/o (CKD+CKE) and w/o (CKD) plateau at 65.81% and 63.44%, respectively. These results highlight the importance of both context and refinement: while refinement alone substantially boosts test executability, the combination with contextual knowledge yields the most effective and consistent improvements.
@@ IV-D RQ3-Model Generality
Interestingly, Qwen2.5-Coder performs on par with GPT-4o-Mini. Under basic prompting, it slightly exceeds GPT-4o-Mini in average coverage. Under our setting, it achieves the highest branch coverage in five out of ten projects and the highest line coverage in four. These results show that with proper enhancement, strong open-source models can match or even surpass closed-source models, offering a cost-effective solution for large-scale automated test generation.
@@ V-A Potential to Reveal Real-World Bugs
JUnitGenie is primarily designed to generate unit tests that achieve high path coverage within focal methods. While it is not explicitly intended for triggering bugs, its ability to explore deeper and more diverse control-flow paths increases the chance of exposing edge cases and latent issues. To assess this potential, we execute all valid tests generated by JUnitGenie across the ten projects. For each failed test execution, we manually examine the expected and actual behaviors to determine if the discrepancy reflects a genuine bug.
@@ V-B Threats to Validity
Finally, results may be influenced by randomness, as EvoSuite and Randoop use randomized search while LLM-based methods rely on probabilistic generation. We mitigate this by fixing the dataset and environment and checking consistency across budgets. Across four time budgets, JUnitGenie consistently achieves the highest branch and line coverage, demonstrating robustness despite stochasticity.

### v2

@@ Abstract
11footnotetext: Xiaoyu Sun is the corresponding author.
@@ I Introduction
To address this, we design and implement JUnitGenie, a path-sensitive framework that guides LLMs in generating high-coverage test cases with minimal cognitive overhead, i.e., by supplying only the essential context needed while filtering out irrelevant information. Specifically, we first build a code knowledge base (CKB) by extracting contextual knowledge (i.e., type information, control flow, and data dependencies) from a Java project under testing. We then traverse each method one control-flow path at a time, retrieve path-specific context from the CKB, and distill it into explicit instructions to guide test generation for that specific path. JUnitGenie applies multi-round refinement to resolve compilation and runtime errors, iteratively generating test cases that are both syntactically correct and executable.
@@
  • We conduct an extensive evaluation to demonstrate the effectiveness of JUnitGenie: (i) it significantly outperforms state-of-the-art tools across real-world Java projects; (ii) ablation studies confirm the contribution of knowledge distillation and test case refinement; (iii) the approach generalizes well across popular foundation models; (iv) it successfully uncovers real bugs in production code.
@@ III-A Context Knowledge Extraction
2) Control-flow information: We then use SootUp [29], which offers precise features for code analysis. In particular, it supports 3-address code intermediate representation Jimple and accurate call-graph construction algorithm Spark [30]. On top of SootUp, JUnitGenie decomposes Java programs into Jimple and decomposes expressions into control flow graph (CFG) paths, modeling method calls as atomic nodes and logical operators as branching edges. Each path has distinct guard conditions, enabling input synthesis for diverse execution scenarios. When extracting paths from a CFG, if a loop is encountered (i.e., a node reappears), we only keep the first iteration to preserve the loop’s logic, and skip subsequent visits within the same path to prevent infinite path extraction.
3) Data-flow dependencies: We again adopt SootUp to extract data-flow information such as variable definitions, field accesses, and dependent method calls to capture how values may influence branching conditions. Specifically, dependent methods are those whose return values affect subsequent control decisions or outputs. For example, as shown in Fig. 1, the return value of contains() (i.e., line 12) determines whether the method proceeds to the following return true statement at line 13. In addition, JUnitGenie also considers dependent variables, including parameters, local variables, and static fields whose values are referenced in conditional statements within the focal method. For instance, the static string variable VOWELS is used in isVowel() to determine whether a character belongs to the set of vowels, thereby influencing the outcome of the conditional branch at line 16. Understanding this context knowledge and their data flow enables LLM to assign appropriate values to variables when constructing test inputs. Together, these dependencies provide a detailed view of how data influences control flow, allowing JUnitGenie to generate inputs that satisfy path-specific conditions and fully exercise the focal method’s behavior.
@@ IV-A Dataset
To evaluate the effectiveness of JUnitGenie, we use real-world Java projects from Defects4J [36], a widely adopted benchmark for automated testing research. We use the latest versions of these projects to reflect modern development practices. In addition, we observe Defects4J includes 17 Java projects with varying JDK requirements. To ensure a consistent and controlled experimental environment, we standardize on JDK 8 as the target platform, given its continued prevalence in open-source libraries and enterprise systems [37, 38, 39]. Based on this criterion, we exclude five projects (JFreeChart, Commons-CLI, Gson, Jsoup, and Math) that require JDK versions beyond 8 and cannot be compiled under our configuration. Additionally, we exclude two projects (Mockito and Closure-Compiler) that are not Maven-based and therefore cannot be built using our toolchain. As a result, 10 Java projects remain and are used as our experimental dataset, as shown in Table I.
Within these 10 projects, not all methods are equally meaningful for evaluating test generation tools. Specifically, methods with trivial logic (e.g., no branching or dependencies) can be easily covered by naive approaches and fail to showcase the strengths of advanced tools. To ensure rigorous evaluation, we focus on methods that are both challenging and representative of real-world complexity. Specifically, we select focal methods that meet two criteria: 1) contain at least one branching condition; and 2) have data dependencies on variables or methods defined outside its own body. These filters yield 8,367 candidate methods across the 10 projects. However, generating tests for all would incur significant computational and token costs. To balance cost with statistical significance, we randomly sample [^2] The sample size is determined based on a confidence level at 95% and a confidence interval at 5 (https://www.surveysystem.com/sscalc.htm). a subset of methods from each project, as summarized in Table I. Hence, the overall dataset contains 2,258 focal methods spread across 762 classes.
@@ IV-B RQ1-Effectiveness
Our first research question concerns the effectiveness of JUnitGenie in test case generation. To this end, we evaluate JUnitGenie by measuring its ability to produce valid tests and comparing its performance with existing SOTA tools. Here, we consider a test case to be valid if (1) the generated code snippet can be successfully compiled and (2) it does not throw an exception before the execution point of the focal method. The first condition ensures that the test case is syntactically correct. The second condition guarantees that the execution context of the focal method under test is properly constructed so that the method can be successfully reached during execution.
@@
In addition, Table II reports the number of valid test cases generated by each tool. Specifically, EvoSuite and Randoop generate 3,232 and 4,018 valid tests, while LLM-based tools ChatTester, CoverUp, HITS, and JUnitGenie produce 3,136, 3,958, 4,756, and 14,190 valid tests, respectively. These results highlight the superior effectiveness of JUnitGenie in generating executable test cases.
Fig. 4(a) further summarizes the distribution of valid test cases per project for each tool. Overall, LLM-based tools tend to produce more valid test cases than heuristic-based approaches. Specifically, the median number of valid test cases per project is 323 for EvoSuite, 415 for Randoop, 260 for ChatTester, 332 for CoverUp, 432 for HITS, and 1,355 for JUnitGenie, with JUnitGenie also exhibiting the highest upper quartile at 1,633. Like most other SOTA approaches, JUnitGenie may fail on focal methods with complex control or data dependencies. Nonetheless, JUnitGenie is still capable of generating more valid test cases than existing tools, demonstrating its effectiveness in leveraging distilled context across diverse codebases.
Moreover, we compare the branch and line coverage achieved by each tool across the 10 Java projects, as shown in Table III. JUnitGenie consistently achieves the highest average coverage, with 56.86% for branch and 61.45% for line coverage, outperforming both heuristic and LLM-based baselines. It delivers particularly strong results on complex projects such as Lang (87.10% branch coverage, 92.41% line coverage) and CSV (76.57% branch coverage, 85.98% line coverage), demonstrating its ability to generate tests that exercise diverse and deep execution paths. Although tools like EvoSuite and Randoop perform well on a few individual projects, their overall average coverage remains substantially lower. Among LLM-based methods, HITS achieves the second-best performance with 32.13% branch and 36.45% line coverage, falling notably behind JUnitGenie. These results further validate the effectiveness of JUnitGenie in producing high-coverage test suites across a range of real-world Java applications. We then break down the comparative results as follows:
@@ IV-C RQ2-Knowledge Acquisition and Utilization
While JUnitGenie demonstrates strong performance in generating high-coverage test cases, the specific contributions of its core components remain unclear. To answer this question, we conduct an ablation study to assess the impact of three key steps: Context Knowledge Extraction (CKE), Context Knowledge Distillation (CKD), and Test Case Refinement (TCR). The study aims to clarify how each component contributes to the overall effectiveness of the test generation process.
Here, we design four ablation baselines to assess the contribution of each component in JUnitGenie. The most basic variant (w/o CKD+CKE+TCR) removes all modules, providing only the focal method and its signature. The second (w/o CKD+CKE) retains only test case refinement, using error messages to repair tests without any contextual input. The third (w/o CKD) includes both context extraction and refinement but omits distillation, presenting raw context without structured guidance. The full JUnitGenie configuration includes all three components. All baselines use GPT-4o-Mini and are evaluated on 2,258 focal methods from 10 real-world Java projects.
TABLE V: Comparison of branch and line coverage using basic prompting (Basic) vs. JUnitGenie (Ours) across four LLMs.
@@
    GPT-4o-Mini CodeLlama DeepSeek-Coder Qwen2.5-Coder     GPT-4o-Mini CodeLlama DeepSeek-Coder Qwen2.5-Coder
@@
Adding test case refinement (w/o CKD+CKE) improves coverage by 12–15%, as it repairs syntactically invalid yet semantically meaningful tests through iterative refinement of compiler error messages. As illustrated by the red line in Fig. 6, refinement raises the executable test rate from 33.49% to 65.81% over five iterations. Other variants show substantial gains over five rounds, with JUnitGenie improving from 45.60% to 69.88%, and w/o (CKD) from 20.32% to 63.44%. These trends highlight the critical role of iterative refinement in recovering test cases, especially when combined with contextual knowledge. It is also worth noting that JUnitGenie, which incorporates both contextual knowledge and refinement, achieves the highest improvement, reaching 69.88% by round five. In comparison, w/o (CKD+CKE) and w/o (CKD) plateau at 65.81% and 63.44%, respectively. These results highlight the importance of both context and refinement: while refinement alone substantially boosts test executability, the combination with contextual knowledge yields the most effective and consistent improvements.
@@ IV-D RQ3-Model Generality
Interestingly, Qwen2.5-Coder performs on par with GPT-4o-Mini. Under basic prompting, it slightly exceeds GPT-4o-Mini in average coverage. Under our setting, it achieves the highest branch coverage in four out of 10 projects and the highest line coverage in three. These results show that with proper enhancement, strong open-source models can match or even surpass closed-source models, offering a cost-effective solution for large-scale automated test generation.
@@ V-A Potential to Reveal Real-World Bugs
JUnitGenie is primarily designed to generate unit tests that achieve high path coverage within focal methods. While it is not explicitly intended for triggering bugs, its ability to explore deeper and more diverse control-flow paths increases the chance of exposing edge cases and latent issues. To assess this potential, we execute all valid tests generated by JUnitGenie across the 10 projects. For each failed test execution, we manually examine the expected and actual behaviors to determine if the discrepancy reflects a genuine bug.
@@ V-B Threats to Validity
Finally, results may be influenced by randomness, as EvoSuite and Randoop use randomized search while LLM-based methods rely on probabilistic generation. We mitigate this by fixing the dataset and environment and checking consistency under four different time budgets. Across all budgets, JUnitGenie consistently achieves the highest branch and line coverage, demonstrating robustness to stochasticity.